Cy=-16/3 when x =-4 and z=3. find y when x=11 and z=-3
Answer by Cromlix(4381)

y varies jointly as x and the inverse of z
y = kx/z
k is the constant of proportionality
To find k:-
y = kx/z
-16/3 = k*(-4)/3
3(-16/3) = -4k
k = -16/-4
k = 4
y = 4x/z
y = 4*11/-3
y = - 44/3
Hope this helps.
:-)